# Improved Prediction Of Ligand-Protein Binding Affinities By Meta-Modeling Ho-Joon Lee1,2,*#, Prashant S. Emani3,*# and Mark B. Gerstein3,4,5,6,7#

### v3

@@ Abstract
The accurate screening of candidate drug ligands against target proteins through computational approaches is of prime interest to drug development efforts. Such virtual screening depends in part on methods to predict the binding affinity between ligands and proteins. Many computational models for binding affinity prediction have been developed, but with varying results across targets. Given that ensembling or meta-modeling methods have shown great promise in reducing model-specific biases, we develop a framework to integrate published force-field-based empirical docking and sequence-based deep learning models. In building this framework, we evaluate many combinations of individual base models, training databases, and several meta-modeling approaches. We show that many of our meta-models significantly improve affinity predictions over base models. Our best meta-models achieve comparable performance to state-of-the-art deep learning tools exclusively based on structures, while allowing for improved database scalability and flexibility through the explicit inclusion of features such as physicochemical properties or molecular descriptors. Overall, we demonstrate that diverse modeling approaches can be ensembled together to gain improvement in binding affinity prediction.
@@ Introduction
Using a benchmark dataset of CASF-2016 7 and a small-scale virtual screening, we demonstrate that our meta-models outperform individual base models and exhibit competitive performance to more sophisticated structure-based DL models, serving as a potentially useful general framework for continual future improvement of ligand-protein binding affinity prediction. The advantages offered by such an approach over existing methods include: (1) scalability by utilizing sequence-based databases for DL models without the need for 3D structures; and (2) flexibility by allowing different base models and physicochemical or molecular properties to be incorporated into meta-models. We believe that such an extensive exploration of these aspects of binding affinity prediction makes a meaningful contribution to the discussion of model generalization.
@@ Methods
We downloaded the PDBbind v2020 refined set from http://www.pdbbind.org.cn/. The dataset consisted of 5,316 target protein-drug ligand complexes. The refined set (labeled RefinedSet) has been filtered by the PDBbind creators from a larger dataset by the removal of complexes that have “obvious problems in 3D structure, binding data or other aspects” 85. Of these, 285 ligand-protein complexes are in the so-called core set (labeled CoreSet), which are high-quality data that are used as the primary test set for the Comparative Assessment of Scoring Functions (CASF) benchmark 7. This CoreSet was separated from the remainder of the PDBbind RefinedSet to serve as an external test set (i.e., benchmark set) for evaluating model performance. Thus, we consider a training set of 5,031 ligand-protein complexes and an external test set of 285 complexes. Binding affinity values we used in this study were the natural logarithms of Kd or Ki values. We did not distinguish Kd and Ki from each other.
@@
We employ the smina implementation 20 of the AutoDock Vina method 19 for the docking component. The smina implementation is designed to make AutoDock easier to use, with larger support for ligand molecular formats and improved minimization algorithms. We used both the default SMINA scoring function and another previously published scoring function, Vinardo 86, provided in smina. We note that the italicized smina describes the program while the capitalized SMINA describes the scoring function. In the meta-models below, we refer to this class of empirical docking methods as E. The tandem use of both scoring functions is meant to address possible idiosyncrasies in the parameterizations of the scoring functions. smina is run as a command-line program using a receptor structure in PDBQT format and a ligand in 3D SDF format. Our in-house Python script includes steps for the preparation of the receptor structures, the creation of the configuration file, and the running of smina (Supporting Information). The output for each scoring function consists of the binding affinities for the 9 (the default option) best docking poses. Seven complexes with multiple ligands (whose names were separated with “&” or “/”) in the PDBbind dataset were not docked as there was some ambiguity as to which ligand to consider in the docking. These complexes were excluded from further analyses.
@@
Previous studies have demonstrated the value of (a) filtering the docking poses output by empirical docking programs based on the degree to which they agree with experimental scores 20 and (b) comparing poses from two or more docking methods, in so-called “consensus” approaches, and filtering based on the degree of agreement to yield a higher quality output 23, 24, 87. Accordingly, we carried out preprocessing of the docking scores by filtering out certain complexes based on the quality of the associated poses assessed by these two criteria. For each complex out of the available 9 docking poses, we calculated the root-mean-square deviation (RMSD) between docking poses in two different ways (described below), and utilized the RMSD value as a cutoff in determining the best set of binding affinities to choose in the final dataset to the meta-models. The goal was to determine if pose quality, either relative to an available experimental gold-standard or relative to the results of another scoring function, would impact both the matching of the predicted binding affinity as well as the downstream prediction process.
@@
An important aspect of these parallel approaches to applying the RMSD filters on the data is that the scores extracted from the empirical docking runs are different when considering the Experimental RMSD filter and the Consensus RMSD filter. This is even the case when considering the unfiltered dataset, where the set of complexes are identical, but the choice of poses used to extract the best docking scores differ between the two RMSD filter approaches. Finally, we note that, in the naming of the model files, we use the shorthand “VvS” to indicate the Consensus RMSD filtering as the comparison involves Vinardo versus SMINA poses.
@@
To understand the contributions of ligand properties to benchmark performance of all the models, we used the web tool SwissADME 88 to extract the properties of the PDBbind CoreSet ligands from their SMILES strings. Also, we investigated the SwissADME features that might be associated with better or worse performance for subsets of complexes by the different groups of tools (docking, DL, and meta-models). To evaluate the contributions of protein features, we extracted the relevant subsets of PDBbind CoreSet PDB IDs from our analyses and fed them into the “Retrieve/ID mapping” tool of the UniProt database 89. We studied the resulting protein properties (e.g. pathway information) for over- or under-representation in our predictions. Details are provided in Supporting Information.
Virtual Screening Benchmarks
@@
To test the possibility of certain tools performing the prediction task with differing degrees of accuracy for certain proteins or even families of proteins, we devised a virtual screening test by selecting 3 most frequently targeted proteins by multiple ligands from the general set of PDBbind v2020 (GeneralSet) excluding the RefinedSet we used for model training. We considered only those proteins in GeneralSet whose sequence similarity to those in RefinedSet was found to be less than 30% by the NCBI Blastp package 90. We also filtered out complexes in the GeneralSet that had only IC50 experimental measurements for consistency with this study. The resulting top 3 proteins are O60341, O15151, and P0C6U8 (UniProt Accession IDs), which are complexed with 20, 8, and 7 ligands, respectively. The resolution of the 35 complexes ranges from 1.33 to 3.50 Angstroms with mean of 2.58 and standard deviation of 0.69.
@@
For a more general applicability, we also performed a virtual screening using an external benchmark dataset, LIT-PCBA, of active and inactive ligands for 15 target proteins 91. Although our models were neither explicitly designed nor trained to distinguish active and inactive ligands from each other (i.e., binary classification for functional activity), our assumption was that active ligands tend to have higher binding affinities (i.e., lower Kd/Ki) for target proteins than inactive ligands. This assumption is weak because there is an unclear causal relationship between binding affinity (Kd, Ki) and functional activity (IC50, EC50). We also note that we used ligand-protein complexes with Kd or Ki from both BindingDB and PDBbind for consistency. LIT-PCBA is a recently developed virtual screening dataset which was reported to be less biased than previous datasets. It is based on PubChem BioAssays which includes 4 potency types of IC50, EC50, Kd, and Ki. We point out that LIT-PCBA did not distinguish the 4 potency types from each other. As a proof-of-concept, with the aforementioned caveat in mind, we randomly selected 10 active and 10 inactive ligands for each of the 15 proteins in LIT-PCBA (i.e., a total of 300 ligand-protein pairs) to predict their binding affinities by our models. The active and inactive ligands were chosen from the combined training and validation sets in LIT-PCBA. We generated three-dimensional SDF files for each of the ligands from the SMILES strings. We also randomly selected only one of the PDB structures (in mol2 format) for each target protein in this virtual screening exercise (further details in Supporting Information). The evaluation was done in two ways: (1) comparisons of predicted binding affinity distributions between active and inactive ligands for each protein with Wilcoxon and Mann-Whitney U tests for significant differences; (2) an enrichment or fraction of active ligands in the top 5 highest-affinity ligands predicted for each protein (i.e., in the top 25% of the 20 evaluated ligands for each protein).
@@ Results
To understand the contributions of the component features, we extracted the feature importance scores for each meta-model (Supplementary Fig. S7). We find that, while the contributions of the docking scores dominate for the EW, ED1, and ED2, these contributions drop to very low values for the other DL-containing meta-models. In all meta-models, SMINA scores contribute more than Vinardo scores. However, given the aforementioned similarity between the SMINA and Vinardo scores, this difference in the contributions between the two docking scores is more likely to be an artifact of the modeling process (and how it deals with nearly collinear features). The other noteworthy aspect of the importance scores is that the performance of ED3, fine-tuned, and PC-based meta-models is dominated by one, or at most two significant features. ED1-F, ED2-F and ED3 are impacted mainly by two DL scores, while the PC models are dominated by PC1. We observe again that PC1 alone is not sufficient to improve the predictions, but requires the assistance of other PCs. The patterns of feature importance scores, however, are not entirely consistent across the different ML meta-models. The top features of the LASSO and Linear Regression meta-models are distinct from the XGBoost top features. The signed feature importance scores in the linear models show less domination of single scores, especially in the PC-based models. In particular, PC1, which is individually highly correlated with experimental binding affinities, is less important in the predictions by the linear meta-models. Given the similar performances of the ML meta-models, it appears that there are several paths to attaining nearly equivalent overall correlation values.
@@
To further parse the differences between the outputs of the prediction tools, we assessed whether certain properties of the proteins or ligands are over- or under-represented in the complexes which were: (1) Best modeled by empirical docking; (2) Best modeled by deep learning; (3) Best modeled by meta-modeling. We also explored the top 50 complexes as predicted by the XGBoost Meta-model for ED-A-P to see if the performance of the meta-models relative to the experimental data was contingent on protein or ligand properties. We extracted protein features from the UniProt database and ligand features from the SwissADME database. We tested the significance of the occurrence of certain annotations (using Fisher’s exact test) and of the distributions of certain quantitative features (using a two-sided Wilcoxon test) for each of the four subsets of complexes. Specifically, we considered whether several UniProt features were annotated more or less frequently, and whether certain Gene Ontology (GO) 93, domain and pathway annotations occurred with different probabilities in the subsets. We also compared the distributions of the ligand SwissADME features. The results (Supplementary Table S9) indicate that there are practically no significant (nominal p-value ) protein features prioritized in the four cases considered. This may be due to the fact that several proteins (and their associated UniProt IDs) are duplicated in the CoreSet. The number of reviewed proteins with UniProtIDs drops from 243 to 60 after removal of all duplicates. The highly significant results with nominal p-values (both Wilcoxon and t-tests) are the distributions of several ligand features including MW, lipophilicity, and solubility measures in the structures for which the empirical docking tools perform the best. There is also some hint of significance for MW, the “#Rotatable bonds” and “#H-bond acceptors” in the structures for which the meta-models perform the best.
@@
Having built our meta-models based on sequence-based DL models, we were interested in comparisons with structure-based tools through voxelization of binding pockets. We focus on the following 3 DL tools by re-training them with our training data for fair comparisons as much as possible as well as one recent tool based on advanced mathematics.
HAC-Net
[…]
FAST
Jones et al. developed meta or fusion models, FAST 42, by combining 3D-CNN and SG-CNN (spatial graph CNN). Due to technical difficulties, we were able to run their pre-trained SG-CNN model trained on the PDBbind-2016 general set excluding the CoreSet. We tested the model against the CoreSet of 266 complexes we used, which resulted in a Pearson correlation of 0.716, worse than HAC-Net. Their MSE is not comparable to ours for the same reason above.
KDeep
KDeep 46 was one of the earlier structure-based DL models with 3D-CNN to predict ligand-protein binding affinity. Their reported Pearson correlation on the PDBbind-2016 CoreSet was 0.82, which we were not able to verify because their code is not open-source. The KDeep model available on the public web-server could be re-trained with limited data size (KdeepTrainer, https://playmolecule.com/KdeepTrainer/). While we were not able to use our full training data due to the limitation, we were able to re-train the KDeep model for 250 epochs with each of 3 training sets of random 1,800 complexes sampled from our training set of the PDBbind-2020 RefinedSet. We then tested the 3 trained models against the CoreSet we used. We obtained Pearson correlations of 0.718–0.766, which are similar to those reported values of 0.701–0.738 by Kwon et al. using the PDBbind-2016 RefinedSet as a training set and 4 different learning rates 50.
PerSpect ML
PerSpect ML 69 is one of the best performing tools in recent years, which reported the highest Pearson correlation of 0.84 in the case of the PDBbind 2016 CoreSet. It is based on topological data analysis and spectral theory for multiscale molecular featurization, which provided a feature set for the gradient boosting tree-based regression. While there was a technical difficulty, due to incomplete code availability, to reproduce or implement PerSpect ML for fair comparisons with the data we used, those recent methods based on advanced mathematics from the Xia group 67-70 showed largely better performances than ours and others. We note, however, that they may generate more than 100,000 features requiring 3D structures or binding poses, which are often unavailable in real-world drug discovery.
Despite those limited comparisons, we conclude that our meta-models or our sequence-based DL models show competitive performances or advantages to more sophisticated structure-based models.
Virtual screening benchmark
@@
To further explore the performance of the models on specific target proteins with multiple ligands, we designed a virtual screening benchmark against 3 target proteins for 35 complexes from the PBDbind v2020 GeneralSet: O60341, O15151, and P0C6U8 (see Methods). The performances of our models for the 3 target proteins are shown in Table 4 in comparison with the 3 tools described above, HAC-Net (re-trained), SG-CNN (FAST; pre-trained), and KDeep (the up-to-date default version on their web server). The pre-trained SG-CNN performed best for O60341 and O15151 in terms of RMSE (2.10 and 1.37, respectively). This was expected because their training data were the PDBbind v2016 GeneralSet, which includes 30 out of the 35 complexes in our benchmark set except 5 complexes for O60341. Otherwise, our meta-models or fine-tuned DL models performed best for all 3 target proteins in terms of both RMSE and Pearson’s correlation. Our best performing models are ED2-F_VvS_101.0_LinReg (meta-model; docking + MW + de novo-trained & fined-tuned DL by linear regression) and D1F|Daylight_AAC (pre-trained & fine-tuned DL model) for O60341 (RMSE = 2.34), EW_VvS_101.0_LinReg (meta-model; docking + MW by linear regression) for O15151 (RMSE = 1.44), and D2F|MPNN-AAC (de novo-trained & fine-tuned DL model) for P0C6U8 (RMSE = 1.01). We consider the Pearson correlation metric less reliable due to the small sample size.
@@
Although the above results suggest that top-ranking high-affinity ligands predicted by our models could be prioritized as hits in virtual screening, we also tested our models using the virtual screening benchmark dataset of LIT-PCBA to distinguish 10 active ligands from 10 inactive ligands for 15 proteins (see Methods; Supplementary Table S10). As our models were not trained for binary classification of active and inactive ligands, our hypothesis was that active ligands tend to have higher affinity scores than inactive ligands. To test the hypothesis, we evaluated the fraction of active ligands in the 5 highest-scoring ligands for each protein (Table 5) and the distribution difference of predicted affinity scores between 10 active and 10 inactive ligands for each protein (Wilcoxon and Mann-Whitney U tests; Supplementary Table S11). The results show that our 5 highest-scoring ligands by either averaged DL models or averaged meta-models tend to be active ligands (i.e., >50%) for 11 out of the 15 proteins and that active ligands tend to have higher affinity scores than inactive ligands for most of the proteins. We also observe that our DL models and meta-models tend to predict higher fractions of active ligands (>=80%) than the baseline docking tools, SMINA or Vinardo (Table 5). Specifically, all the top 5 predicted ligands by D3|Transformer-CNN are active ligands for all the 15 proteins. More than 50% of the top 5 predicted ligands by most of the EW meta-models are active ligands for 13 out of the 15 proteins. We note, however, that, as in the section on prediction synergy, there is variability among the target proteins in terms of the best-performing model. Different protein targets may benefit from the use of different modeling techniques, conditional on their physicochemical properties. This is also observed in terms of the target-specific behavior of the p-values (Supplementary Table S11).

### v4

@@ Abstract
The accurate screening of candidate drug ligands against target proteins through computational approaches is of prime interest to drug development efforts. Such virtual screening depends in part on methods to predict the binding affinity between ligands and proteins. Many computational models for binding affinity prediction have been developed, but with varying results across targets. Given that ensembling or meta-modeling approaches have shown great promise in reducing model-specific biases, we develop a framework to integrate published force-field-based empirical docking and sequence-based deep learning models. In building this framework, we evaluate many combinations of individual base models, training databases, and several meta-modeling approaches. We show that many of our meta-models significantly improve affinity predictions over base models. Our best meta-models achieve comparable performance to state-of-the-art deep learning tools exclusively based on 3D structures, while allowing for improved database scalability and flexibility through the explicit inclusion of features such as physicochemical properties or molecular descriptors. We further demonstrate improved generalization capability by our models using a large-scale benchmark of affinity prediction as well as a virtual screening application benchmark. Overall, we demonstrate that diverse modeling approaches can be ensembled together to gain meaningful improvement in binding affinity prediction.
@@ Introduction
Using 3 diverse benchmark evaluations of CASF-2016 7, the large-scale PDBbind v2020 general set, and a virtual screening dataset of LIT-PCBA 85, we demonstrate that our meta-models outperform individual base models and exhibit competitive performance to more sophisticated structure-based DL models, serving as a potentially useful general framework for continual future improvement of ligand-protein binding affinity prediction. The advantages offered by such an approach over existing methods include: (1) scalability by utilizing sequence-based databases for DL models without the need for 3D structures; and (2) flexibility by allowing different base models and physicochemical or molecular properties to be incorporated into meta-models. We believe that such an extensive exploration of these aspects of binding affinity prediction makes a meaningful contribution to the discussion of model generalization.
@@ Methods
We downloaded the PDBbind v2020 refined set from http://www.pdbbind.org.cn/. The dataset consisted of 5,316 target protein-drug ligand complexes. The refined set (labeled RefinedSet) has been filtered by the PDBbind creators from a larger dataset by the removal of complexes that have “obvious problems in 3D structure, binding data or other aspects” 86. Of these, 285 ligand-protein complexes are in the so-called core set (labeled CoreSet), which are high-quality data that are used as the primary test set for the Comparative Assessment of Scoring Functions (CASF) benchmark 7. This CoreSet was separated from the remainder of the PDBbind RefinedSet to serve as an external test set (i.e., a benchmark set) for evaluating model performance. Thus, we considered a training set of 5,031 ligand-protein complexes and an external test set of 285 complexes. As 266 of 285 affinities were properly predicted upon running the docking tools (see below), we took those 266 complexes as our primary benchmark set. Binding affinity values we used in this study were the natural logarithms of Kd or Ki values. We did not distinguish Kd and Ki from each other.
@@
We employ the smina implementation 20 of the AutoDock Vina method 19 for the docking component. The smina implementation is designed to make AutoDock easier to use, with larger support for ligand molecular formats and improved minimization algorithms. We used both the default SMINA scoring function and another previously published scoring function, Vinardo 87, provided in smina. We note that the italicized smina describes the program while the capitalized SMINA describes the scoring function. In the meta-models below, we refer to this class of empirical docking methods as E. The tandem use of both scoring functions is meant to address possible idiosyncrasies in the parameterizations of the scoring functions. smina is run as a command-line program using a receptor structure in PDBQT format and a ligand in 3D SDF format. Our in-house Python script includes steps for the preparation of the receptor structures, the creation of the configuration file, and the running of smina (Supporting Information). The output for each scoring function consists of the binding affinities for the 9 (the default option) best docking poses. Seven complexes with multiple ligands (whose names were separated with “&” or “/”) in the PDBbind dataset were not docked as there was some ambiguity as to which ligand to consider in the docking. These complexes were excluded from further analyses.
@@
Previous studies have demonstrated the value of (a) filtering the docking poses output by empirical docking programs based on the degree to which they agree with experimental scores 20 and (b) comparing poses from two or more docking methods, in so-called “consensus” approaches, and filtering based on the degree of agreement to yield a higher quality output 23, 24, 88. Accordingly, we carried out preprocessing of the docking scores by filtering out certain complexes based on the quality of the associated poses assessed by these two criteria. For each complex out of the available 9 docking poses, we calculated the root-mean-square deviation (RMSD) between docking poses in two different ways (described below), and utilized the RMSD value as a cutoff in determining the best set of binding affinities to choose in the final dataset to the meta-models. The goal was to determine if pose quality, either relative to an available experimental gold-standard or relative to the results of another scoring function, would impact both the matching of the predicted binding affinity as well as the downstream prediction process.
@@
An important aspect of these parallel approaches to applying the RMSD filters on the data is that the scores extracted from the empirical docking runs are different when considering the Experimental RMSD filter and the Consensus RMSD filter. This is even the case when considering the unfiltered dataset, where the set of complexes are identical, but the choice of poses used to extract the best docking scores differ between the two RMSD filter approaches. Another aspect we note is that experimental structures are not necessarily available for protein-ligand complexes used in many downstream use cases. It is therefore not our intention that users will apply these filters on the prediction dataset. Rather, we utilize the filters on our training data to produce what is hopefully a more informative dataset for the meta-model training. Our evaluations of the relative merits of different filters on the training set are provided in Results. Finally, we note that, in the naming of the model files, we use the shorthand “VvS” to indicate the Consensus RMSD filtering as the comparison involves Vinardo versus SMINA poses.
@@
To understand the contributions of ligand properties to benchmark performance of all the models, we used the web tool SwissADME 89 to extract the properties of the PDBbind CoreSet ligands from their SMILES strings. Also, we investigated the SwissADME features that might be associated with better or worse performance for subsets of complexes by the different groups of tools (docking, DL, and meta-models). To evaluate the contributions of protein features, we extracted the relevant subsets of PDBbind CoreSet PDB IDs from our analyses and fed them into the “Retrieve/ID mapping” tool of the UniProt database 90. We studied the resulting protein properties (e.g. pathway information) for over- or under-representation in our predictions. Details are provided in Supporting Information.
Comparison to other tools
[…]
GeneralSet Benchmark
To assess model generalizability on a large scale, we performed an external validation of binding affinities for complexes in the PDBbind v2020 GeneralSet as another benchmark in addition to the CoreSet. The full GeneralSet consists of 19,443 complexes and is not as carefully curated as the RefinedSet. Hence, we filtered complexes according to the following rules:
1.
Filter out all complexes belonging to the RefinedSet and CoreSet.
2.
Remove all complexes assayed by NMR.
3.
Remove all complexes with an IC50 dissociation measure (as opposed to Kd or Ki).
4.
Only include complexes whose PDB structures were published after the year 2000.
5.
Finally, not all complexes had successful docking calculations for both Vinardo and SMINA. We left out complexes with any predicted binding affinity of zero.
The resulting subset included 5,261 complexes, which is larger than our training set of 5,031 complexes from the RefinedSet. Therefore, it is a more challenging benchmark than the CoreSet. We ran the docking and our model predictions using the same settings as for the CoreSet.
Ranking Benchmark
@@
To test the possibility of certain tools performing the prediction task with differing degrees of accuracy for certain proteins or even families of proteins, we devised a virtual screening test by selecting 3 most frequently targeted proteins by multiple ligands from the PDBbind GeneralSet excluding the RefinedSet we used for model training. We considered only those proteins in GeneralSet whose sequence similarity to those in RefinedSet was found to be less than 30% by the NCBI Blastp package 92. We also filtered out complexes in the GeneralSet that had only IC50 experimental measurements for consistency with this study. The resulting top 3 proteins are O60341, O15151, and P0C6U8 (UniProt Accession IDs), which are complexed with 20, 8, and 7 ligands, respectively. The resolution of the 35 complexes ranges from 1.33 to 3.50 Angstroms with mean of 2.58 and standard deviation of 0.69.
Virtual Screening
@@
For a more general applicability, we also performed a virtual screening using an external benchmark dataset, LIT-PCBA, of active and inactive ligands for 15 target proteins 85. Although our models were neither explicitly designed nor trained to distinguish active and inactive ligands from each other (i.e., binary classification for functional activity), our assumption was that active ligands tend to have higher binding affinities (i.e., lower Kd/Ki) for target proteins than inactive ligands. This assumption is weak because there is an unclear causal relationship between binding affinity (Kd, Ki) and functional activity (IC50, EC50). We also note that we used ligand-protein complexes with Kd or Ki from both BindingDB and PDBbind for consistency. LIT-PCBA is a recently developed virtual screening dataset which was reported to be less biased than previous datasets. It is based on PubChem BioAssays which includes 4 potency types of IC50, EC50, Kd, and Ki. We point out that LIT-PCBA did not distinguish the 4 potency types from each other. As a proof-of-concept, with the aforementioned caveat in mind, we randomly selected active and inactive ligands maintaining the ratio 1:2 (number of active ligands:number of inactive ligands) for each of the 15 proteins in LIT-PCBA to predict their binding affinities by our models. Specifically, if the number of active ligands was < 100, we included them all; if greater than or equal to 100, we randomly sampled 100 active ligands. The inactive ligand numbers were then selected to maintain the aforementioned 1:2 ratio. This gave us a total of 1,050 active ligand-target protein pairs and 2,100 inactive ligand-target protein pairs. The active and inactive ligands were chosen from the combined training and validation sets in LIT-PCBA. We generated three-dimensional SDF files for each of the ligands from the SMILES strings. We also randomly selected only one of the PDB structures (in mol2 format) for each target protein in this virtual screening exercise (further details in Supporting Information). The evaluation was done in 4 ways: (1) comparisons of predicted binding affinity distributions between active and inactive ligands for each protein with Welch’s t-test and the Mann-Whitney U test for significant differences; (2) top5 recall: an enrichment or fraction of active ligands (hits) in the top 5 highest-affinity ligands predicted for each protein; (3) similarly, top10 recall for hits in the top 10 ligands; and (4) precision for hits in the number of the top ligands equal to that of all active ligands for each protein.
@@ Results
To understand the contributions of the component features, we extracted the feature importance scores for each meta-model (Supplementary Figure S7). We find that, while the contributions of the docking scores dominate for the EW, ED1, and ED2, these contributions drop to very low values for the other DL-containing meta-models. In all meta-models, SMINA scores contribute more than Vinardo scores. However, given the aforementioned similarity between the SMINA and Vinardo scores, this difference in the contributions between the two docking scores is more likely to be an artifact of the modeling process (and how it deals with nearly collinear features). The other noteworthy aspect of the importance scores is that the performance of ED3, fine-tuned, and PC-based meta-models is dominated by one, or at most two significant features. ED1-F, ED2-F and ED3 are impacted mainly by two DL scores, while the PC models are dominated by PC1. We observe again that PC1 alone is not sufficient to improve the predictions, but requires the assistance of other PCs. The patterns of feature importance scores, however, are not entirely consistent across the different ML meta-models. The top features of the LASSO and Linear Regression meta-models are distinct from the XGBoost top features. The signed feature importance scores in the linear models show less domination of single scores, especially in the PC-based models. In particular, PC1, which is individually highly correlated with experimental binding affinities, is less important in the predictions by the linear meta-models. Given the similar performances of the ML meta-models, it appears that there are several paths to attaining nearly equivalent overall correlation values.
@@
To further parse the differences between the outputs of the prediction tools, we assessed whether certain properties of the proteins or ligands are over- or under-represented in the complexes which were: (1) Best modeled by empirical docking; (2) Best modeled by deep learning; (3) Best modeled by meta-modeling. We also explored the top 50 complexes as predicted by the XGBoost Meta-model for ED-A-P to see if the performance of the meta-models relative to the experimental data was contingent on protein or ligand properties. We extracted protein features from the UniProt database and ligand features from the SwissADME database. We tested the significance of the occurrence of certain annotations (using Fisher’s exact test) and of the distributions of certain quantitative features (using a two-sided Wilcoxon test) for each of the four subsets of complexes. Specifically, we considered whether several UniProt features were annotated more or less frequently, and whether certain Gene Ontology (GO) 94, domain and pathway annotations occurred with different probabilities in the subsets. We also compared the distributions of the ligand SwissADME features. The results (Supplementary Table S9) indicate that there are practically no significant (nominal p-value ) protein features prioritized in the four cases considered. This may be due to the fact that several proteins (and their associated UniProt IDs) are duplicated in the CoreSet. The number of reviewed proteins with UniProtIDs drops from 243 to 60 after removal of all duplicates. The highly significant results with nominal p-values (both Wilcoxon and t-tests) are the distributions of several ligand features including MW, lipophilicity, and solubility measures in the structures for which the empirical docking tools perform the best. There is also some hint of significance for MW, the “#Rotatable bonds” and “#H-bond acceptors” in the structures for which the meta-models perform the best.
@@
Having built our meta-models based on sequence-based DL models, we were interested in comparisons with structure-based tools through voxelization of binding pockets. As described in Methods, we focused on the following 3 DL tools by re-training them with our training data for fair comparisons as much as possible as well as one recent tool based on advanced mathematics. HAC-Net 47 was trained on the PDBbind-2020 general set of >18,800 ligand-protein complexes. Their best performance was a Pearson correlation of 0.846 for the PDBbind-2016 CoreSet of 290 complexes. We note that we focused on the PDBbind-2020 RefinedSet rather than the general set because of two reasons: (1) higher quality and (2) limited resources. The re-trained HAC-Net with our training set showed a Pearson correlation of 0.756 for the CoreSet we used, which is worse than our best meta-model performance of 0.777. Their MSE values are not comparable to ours because they used -log10(Kd/Ki) while we used ln(Kd/Ki), which show different distributions of binding affinities for model training. FAST 42 was trained on the PDBbind-2016 general set excluding the CoreSet. We tested the pre-trained SG-CNN of FAST against the CoreSet we used, which resulted in a Pearson correlation of 0.716, worse than HAC-Net. KDeep 46 was one of the earlier structure-based 3D-CNN models. Their reported Pearson correlation on the PDBbind-2016 CoreSet was 0.82, which we were not able to verify because their code is not open-source. With 3 re-trained KDeep models using 3 limited training sets from KdeepTrainer (see Methods), We obtained Pearson correlations of 0.718–0.766 for the CoreSet we used, which are similar to those reported values of 0.701–0.738 by Kwon et al. using the PDBbind-2016 RefinedSet as a training set and 4 different learning rates 50. PerSpect ML 69 recently reported the highest Pearson correlation of 0.84 in the case of the PDBbind 2016 CoreSet. While we were not able to confirm their result due to incomplete code availability, those recent methods based on advanced mathematics from the Xia group 67-70 showed largely better performances than ours and others. We note, however, that they may generate more than 100,000 features requiring 3D structures or binding poses, which are often unavailable in real-world drug discovery. Despite these limited comparisons, we conclude that our meta-models and our sequence-based DL models show competitive performances or advantages to more sophisticated structure-based models.
GeneralSet benchmark
In addition to the CASF-2016 benchmark, we tested our meta-models against the PBDbind2020 GeneralSet of 5,261 complexes as a benchmark (Methods; Table 4). Although the size of the GeneralSet benchmark is as large as our training data, we observe performance improvements of our meta-models over the individual base models (Supplementary Figure S8). As expected, the performance for this benchmark is worse than that for CASF-2016 (PCC <0.60 vs. >0.76) because of its lower data quality and the data size being similar to the training data. As for the structure-based tools mentioned above, Pearson’s correlation by the re-trained HAC-Net was 0.001 (excluding 22 complexes that failed in pre-processing) and that by KDeep (the web-server default model of ACEBind2023-GraphNet as of Sept. 2, 2024) was 0.318 (excluding 71 complexes that failed in pre-processing), which performed worse than our meta-models or DL models and similar to the docking tools. We note that the huge RMSE by HAC-Net is due to very large prediction errors for more than 1,000 complexes, including 505 complexes with positive predictions. The PDBbind2016-GeneralSet pre-trained SG-CNN of FAST achieved Pearson’s correlation of 0.526 (excluding 124 complexes that failed in pre-processing), which is also a worse performance than our meta-models and DL models. Although their RMSE is better than ours, it is not a fair metric for comparison as explained in the previous section. Also, we note that 3,314 out of 5,261 complexes in our benchmark set are part of the training set of PDBbind2016 GeneralSet by FAST.
Ranking benchmark
@@
To further explore the performance of the models on specific target proteins with multiple ligands, we designed a binding affinity ranking benchmark against 3 target proteins for 35 complexes from the PBDbind v2020 GeneralSet: O60341, O15151, and P0C6U8 (see Methods). The performances of our models for the 3 target proteins are shown in Table 5 in comparison with the 3 tools described above, HAC-Net (re-trained), SG-CNN (FAST; pre-trained), and KDeep (the up-to-date default version on their web server). The pre-trained SG-CNN performed best for O60341 and O15151 in terms of RMSE (2.10 and 1.37, respectively). This was expected because their training data were the PDBbind v2016 GeneralSet, which includes 30 out of the 35 complexes in our benchmark set except 5 complexes for O60341. Otherwise, our meta-models or fine-tuned DL models performed best for all 3 target proteins in terms of both RMSE and Pearson’s correlation. Our best performing models are ED2-F_VvS_101.0_LinReg (meta-model; docking + MW + de novo-trained & fined-tuned DL by linear regression) and D1F|Daylight_AAC (pre-trained & fine-tuned DL model) for O60341 (RMSE = 2.34), EW_VvS_101.0_LinReg (meta-model; docking + MW by linear regression) for O15151 (RMSE = 1.44), and D2F|MPNN-AAC (de novo-trained & fine-tuned DL model) for P0C6U8 (RMSE = 1.01). We consider the Pearson correlation metric less reliable due to the small sample size.
@@
Application to virtual screening
Although the above results suggest that top-ranking high-affinity ligands predicted by our models could be prioritized as hits in virtual screening, we also tested our models using the virtual screening benchmark dataset of LIT-PCBA to distinguish active ligands from inactive ligands for 15 proteins (see Methods; Supplementary Table S10). As our models were not trained for binary classification of active and inactive ligands, our hypothesis was that active ligands tend to have higher affinity scores than inactive ligands. To test the hypothesis, we evaluated our models and the 2 docking tools using 3 performance metrics (top5 recall, top10 recall, and precision) for each protein (Methods; Table 6) and the distribution difference of predicted affinity scores between all active and all inactive ligands for each protein (Welch’s t-test and the Mann-Whitney U test; Supplementary Table S11). The 3 performance metrics show that our averaged DL models and meta-models tend to be better at predicting active ligands than the 2 docking tools (i.e., >50% in blue in Table 6). The docking tools performed better for 3 proteins, but their performance values are mostly 20% or less otherwise (in red in Table 6). As expected, active ligands tend to have higher affinity scores than inactive ligands for most of the proteins individually, and when aggregated across all proteins (Supplementary Table S11). We emphasize that, as in the section on prediction synergy, there is variability among the target proteins in terms of the best-performing model. Different protein targets may benefit from the use of different modeling techniques, conditional on their physicochemical properties. This is also observed in terms of the target-specific behavior of the p-values (Supplementary Table S11). All those structure-based models mentioned above, HAC-Net, FAST, KDeep, and the models from the Xia group (e.g. PerSpect ML), require input data of docked poses and hence cannot be used in the LIT-PCBA virtual screening benchmark or are not generally applicable to identification of active vs. inactive ligands.